OpenStudy (knockknock):

Consider the function represented by the equation 1/2j+1/4k=3.Which shows the equation written in function notation with j as the independent variable

OpenStudy (anonymous):

If j is the independent variable, you need to solve the equation for k.

OpenStudy (anonymous):

Is it just the 2 and 4 in the denominator, or are the j and k there too?

OpenStudy (knockknock):

I can give you the 4 options for it. F(j)=-2j+12 F(j)=-1/2j+6 F(k)=-2k+12 F(k)=-1/2k+6

OpenStudy (anonymous):

No, I meant what is the original equation? Which of these?\[\frac{ 1 }{ 2 }j+\frac{ 1 }{ 4 }k=3\]\[\frac{ 1 }{ 2j }+\frac{ 1 }{ 4k }=3\]

OpenStudy (knockknock):

My mistake it's the first one that you have (I can't write my equations like that for some reason)

OpenStudy (knockknock):

I just want to know how to start solving it so I can find out the answer

OpenStudy (anonymous):

Multiply by 4 to clear the fractions

OpenStudy (anonymous):

because it's the least common denominator

OpenStudy (anonymous):